# Claas Voelcker

### v1

@@ Abstract
The idea of decision-aware model learning, that models should be accurate where it matters for decision-making, has gained prominence in model-based reinforcement learning. While promising theoretical results have been established, the empirical performance of algorithms leveraging a decision-aware loss has been lacking, especially in continuous control problems. In this paper, we present a study on the necessary components for decision-aware reinforcement learning models and we showcase design choices that enable well-performing algorithms. To this end, we provide a theoretical and empirical investigation into prominent algorithmic ideas in the field. We highlight that empirical design decisions established in the MuZero line of works are vital to achieving good performance for related algorithms, and we showcase differences in behavior between different instantiations of value-aware algorithms in stochastic environments. Using these insights, we propose the Latent Model-Based Decision-Aware Actor-Critic framework ($\lambda$-AC) for decision-aware model-based reinforcement learning in continuous state-spaces and highlight important design choices in different environments.
@@ 1 Introduction
In model-based reinforcement learning, an agent collects information in an environment and uses it to learn a model of the world. This model helps to accelerate and improve the value function estimation or the agent’s policy [65, 12, 41, 29, 50, 61, 5, 30]. However, as environment complexity increases, learning a model becomes a more and more challenging task. The resulting model errors can impact the learned policy [60, 36, 59, 67, 46, 33, 39] leading to worse performance. In many scenarios, agents are faced with a "big world" setting where learning to accurately simulate the entire environment is infeasible. When faced with such complex environments, it is necessary to decide what to model. Otherwise, capacity would be spent on irrelevant aspects, e.g., modelling the evolution of clouds in a self-driving vehicle’s video feed.
@@
Recently, the paradigm of decision-aware model learning (DAML) [17, 16, 1, 70] or value equivalence (VE) [25, 26] has been proposed. The core idea is that a learned model should make predictions that result in correct value function estimation. Two prominent algorithms in the space of decision-aware model learning are the IterVAML [16] and MuZero [61] algorithms. While IterVAML is an algorithm with clear theoretical guarantees, several difficulties in adapting the loss for well-performing empirical implementations have been highlighted in the literature [44, 70]. MuZero, on the other hand, has been shown to perform well in discrete control tasks [61, 71] but has received little theoretical investigation. Thus, understanding the role of different value-aware losses and determining the factors for achieving strong performance is an open research problem.
In this work, we address this problem by establishing similarities and differences between these two approaches. Our goal is building a well-performing algorithmic framework and providing insight into the relative strengths and weaknesses of each. We focus on the continuous control setting as previous work has mostly focused on the discrete case.
@@
The main differences between previous works trying to build on the IterVAML and MuZero algorithms are neural network architecture design decisions. MuZero is built on value prediction networks, which explicitly incorporate neural networks and a latent world model, while IterVAML is designed for arbitrary model choices. We showcase that using latent models is indeed crucial for achieving good performance with both decision-aware losses.
@@
As a novel theoretical contribution, we prove that a unique feature of IterVAML-based models is that they can achieve low error in stochastic environments, ven when using deterministic world models. The MuZero loss however results in a bias in stochastic environments, while IterVAML provides an unbiased loss function. We show that this bias leads to a quantifiable difference in performance.
Finally, we show when decision-aware losses provide benefits in practice and propose a novel way of combining the value function-oriented algorithms with policy gradient estimation approaches. We propose the Latent Model-Based Decision-Aware Actor-Critic ($\lambda$-AC) framework that combines three core components: a latent world model, a value-aware model loss function, and a model-based actor-critic algorithm for reinforcement learning.
@@
In summary, our contributions are a) showing that IterVAML is a stable loss when a latent model is used, b) proving and verifying empirically that the MuZero loss is biased in stochastic environments, and c) establishing the $\lambda$-AC framework for decision-aware model-based RL.
@@ 2 Background
We consider a standard Markov decision process (MDP) [57, 66] $(\mathcal{X},\mathcal{A},p,r,\gamma)$, with state space $\mathcal{X}$, action space $\mathcal{A}$, transition kernel $p(x^{\prime}|x,a)$, reward function $r:\mathcal{X}\times\mathcal{A}\xrightarrow[]{}\mathbb{R}$, and discount factor $\gamma\in[0,1)$. The goal of an agent is to optimize the obtained average discounted infinite horizon reward under its policy: $\max_{\pi}\mathbb{E}_{\pi,p}\left[\sum_{t=0}^{\infty}\gamma^{t}r(x_{t},a_{t})\right]$.
Given a policy $\pi(a|x)$, the value function is defined as the expected return conditioned on a starting state $x$: $V^{\pi}(x)=\mathbb{E}_{\pi}\left[\sum_{t\geq 0}\gamma^{t}r_{t}|x_{0}=x\right]$, where $r_{t}=r(x_{t},a_{t})$ is the reward at time $t$. The value function is the unique stationary point of the Bellman operator: $\mathcal{T}_{p}(V^{\pi})(x)=\mathbb{E}_{\pi(a|s)}\left[r(x,a)+\gamma\mathbb{E}_{p(x^{\prime}|x,a)}\left[V^{\pi}(x^{\prime})\right]\right]$. The action-value function is defined similarly as $Q^{\pi}(x,a)=\mathbb{E}_{\pi}\left[\sum_{t\geq 0}\gamma^{t}r_{t}|x_{0}=x,a_{0}=a\right]$. In this work, we do not assume access to the reward functions and denote learned approximation as $\hat{r}(x,a)$.
An environment model is a function that approximates the behavior of the transition kernel of the ground-truth MDP. [^1] In this paper, we will generally use the term model to refer to an environment model, not to a neural network, to keep consistent with the reinforcement learning nomenclature. Learned models are used to augment the reinforcement learning process in several ways [65, 9, 33, 29, 5], and depending on the way they are used and trained the behavior and guarantees of the RL agent change significantly.
We will use $\hat{p}(x^{\prime}|x,a)$ to refer to approximate model, and use $\hat{f}(x,a)$ when referring specifically to deterministic models. All models which take a state $x\in\mathcal{X}$ as an input and map to distributions over $\mathcal{X}$ will be referred to as observation-space models. An alternative are latent models of the form $\hat{p}(z^{\prime}|z,a)$, where $z\in\mathcal{Z}$ is a representation of a state $x\in\mathcal{X}$ given by $\phi:\mathcal{X}\rightarrow\mathcal{Z}$. We present further details on latent model learning in section 3.
The notation $\hat{x}^{(n)}_{i}$ refers to a $n$-step rollout obtained from a model by starting in state $x_{i}$ and sampling the model for $n$-steps. We will use $\hat{p}^{n}\left(x^{(n)}\middle|x\right)$ and $\hat{f}^{n}(x)$ to refer to the $n$-th step model prediction.
@@ 2.1 Decision-Aware Model Losses
The losses of the decision-aware learning framework share the goal of finding models that provide good value function estimates. Instead of simply learning a model using maximum likelihood estimation, the value function is directly used in the model update.
Given a distribution $\mu$ over the state space and a value function approximation $\hat{V}$, the IterVAML loss [16] is defined as
$\mathcal{L}_{\text{IterVAML}}(\hat{p};\mu,\hat{V},p)=\mathbb{E}_{(x,a)\sim\mu}\left[\left|\int_{\mathcal{X}}\left(\hat{p}(x^{\prime}|x,a)-p(x^{\prime}|x,a)\right)\hat{V}(x^{\prime})\mathrm{d}x^{\prime}\right|^{2}\right].$ (1)
Although not originally presented in [16], this formulation can be easily extended to a multi-step version $\mathcal{L}^{n}_{\text{IterVAML}}$. In practice, we assume access to a dataset of state and reward sequences $\mathcal{D}=\{x_{i_{1}},r_{i_{1}},\dots,x_{i_{n}},r_{i_{n}}\}_{i=1}^{N}$ with $x_{i_{0}}$ sampled from $\mu$ and subsequent states sampled according to the transition function. Note that the index $1$ here denotes the start of a sequence, not the start of an episode, and sequences can be overlapping. This is similar to other losses in related works [29, 5, 23]. Actions are dropped for legibility. Then, the sample-based IterVAML loss is
^$\hat{\mathcal{L}}^{n}_{\text{IterVAML}}(\hat{p};\hat{V},\mathcal{D})$ $=\frac{1}{N\cdot n}\sum_{i=1}^{N}\sum_{j=1}^{n}\left[\mathbb{E}_{\hat{x}^{(j)}\sim\hat{p}^{j}(\cdot|x_{i_{1}})}\left[\hat{V}\left(\hat{x}^{(j)}\right)\right]-\hat{V}\left(x_{i_{j}}\right)\right]^{2}.$ (2)
The population-based MuZero loss is not defined in the literature, but the sample-based version is. It uses the stop gradient operation popularized by Mnih et al. [47], which we denote with $(\cdot)_{\text{sg}}$. Similar to Mnih et al. [47], it also uses a target network for the value function, which we denote as $V_{\text{target}}$. The resulting MuZero loss is given as
^$\hat{\mathcal{L}}^{n}_{\text{MuZero}}(\hat{f},\hat{V};\mathcal{D})$ $=\frac{1}{N\cdot n}\sum_{i=1}^{N}\sum_{j=i}^{n}\left[\hat{V}\left(\hat{f}^{j}(x_{i_{1}})\right)-\left[r_{i_{j}}+\gamma V_{\text{target}}\left(x_{i_{j+1}}\right)\right]_{\text{sg}}\right]^{2}.$ (3)
Note that the real environment reward and next states are used in the $j$-th step. The main difference between the two loss formulations is that IterVAML uses a squared minimization between two value function outputs, while MuZero computes the value function target via a bootstrap approximation. For the $1$-step case, this is illustrated in Figure 1. The MuZero loss is also used for updating both the value function and the model, while IterVAML is only designed to update the model.
@@
2.2 Model-based value function learning in continuous control
Both IterVAML and MuZero can use the model to approximate the value function target. In its original formulation, MuZero is used with a Monte Carlo Tree Search procedure to expand the model which is not directly applicable in continuous state spaces. A simple alternative for obtaining a value function estimate from the model in continuous state spaces is known as model value expansion (MVE) [18]. In MVE, short horizon roll-outs of model predictions $[{\hat{x}^{0},\dots,\hat{x}^{n}}]$ are computed using the current policy estimate, with a real sample from the environment as a starting point $x^{(0)}$. Using these, n-step value function targets are computed. These targets represent a bootstrap estimate of the value function starting from the j-th step of the model rollout resulting in
$Q_{\text{target}}^{j}\left(x^{(0)}\right)=\sum_{i=j}^{n-1}\gamma^{i-j}\hat{r}(\hat{x}^{(i+j)})+\gamma^{n-j}\bar{Q}\left(\hat{x}^{(n)},\pi\left(\hat{x}^{(n)}\right)\right).$ (4)
The value function estimate $\hat{V}$ or state-action value estimate $\hat{Q}$ is then updated using Fitted Q-Iteration [14]. No gradients are passed to the value function target, analogous to methods such as DQN [47] only the estimate $\hat{Q}$ is updated. In the experiment, the implementation is based on TD3 [21].
The $j$-step formulation ensures that intermediate value functions at states which only appear in the model rollout are also updated based on the approximate value function of the model predictions. Note that the value function gradient is not passed into the model, a core difference between this approach and the MuZero algorithm outlined above.
2.3 Model-based policy updates
In continuous control tasks, estimating a policy is a crucial step for effective learning algorithms. In this work, we use model-based value gradients [31, 5] and as an ablation deterministic policy gradients [63, 43]. Model-based policy gradients [31] are obtained by differentiating the full n-step MVE target (Equation 4) with regard to the policy. If differentiable policies and models are used, a policy gradient over the complete computational graph of the model rollout is obtained. In the case of stochastic environments or stochastic policies, this approach only provides a biased approximation of the true policy gradient [45]. However differentiating the model greatly reduces the variance of the policy gradient estimates and is therefore useful nonetheless. We evaluate both variants in section 5.
3 The importance of latent spaces for applied decision-aware model learning
Figure 2: Comparison on the use of an explicit latent space with different loss functions. All models profit immensely from the addition of an explicit latent space transformation. Dashed lines represent the mean of the latent space result.
One of the differences between previous work on MuZero and IterVAML is that the former was established using a latent model design, while the original IterVAML algorithm is presented without establishing a model implementation. However, when directly applied on top of a state-space prediction model, IterVAML has been shown to diverge, leading to several works claiming that it is an impractical algorithm [44, 70]. There is no a priori reason as to why a latent model should not be used with IterVAML, or why MuZero has to be used with one. We hypothesize that this design difference is the main reasons for the differing empirical performance of IterVAML and MuZero in experiments.
The loss functions presented in subsection 2.1 do not change significantly due to the introduction of the model structure. However, there are two core differences that appear due to incorporating latent embeddings into the model loss.
Firstly, the geometry of the input to the value function changes. As sharp value function gradients have been hypothesized to be a major concern for IterVAML [70], it stands to reason that introducing the ability to learn a representation space in the model through a latent embedding alleviates this issue. Theoretical backing for this hypothesis has been provided in Gelada et al. [22] and Tang et al. [68].
@@ 3 The importance of latent spaces for applied decision-aware model learning
Furthermore, the MuZero loss can be insufficient for representation learning on its own [71]. Its performance is improved by using an auxiliary latent self-prediction loss based on the Bootstrap your own latent (BYOL) loss [24]. Several different versions of this loss have been proposed for RL [22, 62, 68], we consider a simple $n$-step variant:
@@
^$\hat{\mathcal{L}}^{n}_{\text{latent}}\left(\hat{f},\phi;\mathcal{D}\right)=\frac{1}{N\cdot n}\sum_{i=1}^{N}\sum_{j=1}^{n}\left[\hat{f}^{(j)}(\phi(x_{i_{1}}))-\left[\phi(x_{i_{j}})\right]_{\text{sg}}\right]^{2}.$ (5)
To test the impact of both the latent space formulation and the auxiliary latent self-prediction loss, we evaluated both MuZero and IterVAML as model loss functions with and without the stabilizing BYOL loss. For clarity, pseudocode is found in subsection B.2. The experiments are repeated over eight random seeds and shaded areas mark three $\sigma$ of the standard error of the mean over seeds.
As is evident from Figure 2 the latent space is vital to achieve good performance, even in the relatively simple Walker-run environment. Without the stabilization loss, we were unable to prevent divergence with IterVAML, similar to the results presented by Voelcker et al. [70].
This highlights that previous negative results on the efficacy of IterVAML are largely due to suboptimal choices in the model implementation, not due to fundamental limitations of the algorithm. In addition, the stabilizing loss is essential for IterVAML, while it can be dropped in MuZero [71]. We conjecture that MuZero is more stable without auxiliary losses due to incorporating the ground truth reward into its joint model-value function loss, an advantage that IterVAML does not have.

### v3

@@ Abstract
The idea of decision-aware model learning, that models should be accurate where it matters for decision-making, has gained prominence in model-based reinforcement learning. While promising theoretical results have been established, the empirical performance of algorithms leveraging a decision-aware loss has been lacking, especially in continuous control problems. In this paper, we present a study on the necessary components for decision-aware reinforcement learning models and we showcase design choices that enable well-performing algorithms. To this end, we provide a theoretical and empirical investigation into algorithmic ideas in the field. We highlight that empirical design decisions established in the MuZero line of works, most importantly the use of a latent model, are vital to achieving good performance for related algorithms. Furthermore, we show that the MuZero loss function is biased in stochastic environments and establish that this bias has practical consequences. Building on these findings, we present an overview of which decision-aware loss functions are best used in what empirical scenarios, providing actionable insights to practitioners in the field.
@@ 1 Introduction
In model-based reinforcement learning, an agent collects information in an environment and uses it to learn a model of the world to accelerate and improve value estimation or the agent’s policy (Sutton, 1990; Deisenroth & Rasmussen, 2011; Hafner et al., 2020; Schrittwieser et al., 2020). However, as environment complexity increases, learning a model becomes more and more challenging. These model errors can impact the learned policy (Schneider, 1997; Kearns & Singh, 2002; Talvitie, 2017; Lambert et al., 2020) leading to worse performance. In many realistic scenarios, agents are faced with a "big world" where learning to accurately simulate the environment is infeasible. When faced with complex environments, deciding what aspects of the environment to model is crucial. Otherwise, capacity would be spent on irrelevant aspects, e.g., modelling clouds in a vehicle’s video feed.
@@
Recently, the paradigm of decision-aware model learning (DAML) (Farahmand et al., 2017) or value equivalence (Grimm et al., 2020; 2021) has been proposed. The core idea is that a model should make predictions that result in correct value estimation. The two most prominent approaches in the space of decision-aware model learning are the IterVAML (Farahmand, 2018) and MuZero (Schrittwieser et al., 2020) algorithms. While IterVAML is a theoretically grounded algorithm, difficulties in adapting the loss for empirical implementations have been highlighted in the literature (Lovatto et al., 2020; Voelcker et al., 2022). MuZero on the other hand has been shown to perform well in discrete control tasks (Schrittwieser et al., 2020; Ye et al., 2021) but has received little theoretical investigation. Thus, understanding the role of different value-aware losses and determining the factors for achieving strong performance is an open research problem.
Research question:
In this work, we ask three question: (a) What design decisions explain the performance difference between IterVAML and Muzero? (b) What are theoretical or practical differences between the two? (c) In what scenarios do decision-aware losses help in model-based reinforcement learning? Experimental evaluations focus on the state-based DMC environments (Tunyasuvunakool et al., 2020), as the focus of our paper is not to scale up to image-based observations.
Contributions:
@@
The main differences between previous works trying to build on the IterVAML and MuZero algorithms are neural network architecture design and the value function learning scheme. MuZero is built on value prediction networks (Oh et al., 2017), which explicitly incorporate a latent world model, while IterVAML is designed for arbitrary model choices. We show that the use of latent models can explain many previously reported performance differences between MuZero and IterVAML (Voelcker et al., 2022; Lovatto et al., 2020).
@@
As a theoretical contribution, we first show that IterVAML and MuZero-based models can achieve low error in stochastic environments, even when using deterministic world models, a common empirical design decision (Oh et al., 2017; Schrittwieser et al., 2020; Hansen et al., 2022). To the best of our knowledge, we are the first to prove this conjecture and highlight that it is a unique feature of the IterVAML and MuZero losses. The MuZero value function learning scheme however results in a bias in stochastic environments. We show that this bias leads to a quantifiable difference in performance. Finally, we show that the model learning component of MuZero is similar to VAML, which was previously also noted by Grimm et al. (2021).
@@
In summary, our contributions are a) showing that IterVAML is a stable loss when a latent model is used, b) proving and verifying that the MuZero value loss is biased in stochastic environments, and c) showing how these algorithms provide benefit over decision-agnostic baseline losses.
@@ 2 Background
Reinforcement Learning:
We consider a standard Markov decision process (MDP) (Puterman, 1994; Sutton & Barto, 2018) $(\mathcal{X},\mathcal{A},p,r,\gamma)$, with state space $\mathcal{X}$, action space $\mathcal{A}$, transition kernel $p(x^{\prime}|x,a)$, reward function $r:\mathcal{X}\times\mathcal{A}\xrightarrow[]{}\mathbb{R}$, and discount factor $\gamma\in[0,1)$. The goal of an agent is to optimize the obtained average discounted infinite horizon reward under its policy: $\max_{\pi}\mathbb{E}_{\pi,p}\left[\sum_{t=0}^{\infty}\gamma^{t}r(x_{t},a_{t})\right]$.
Given a policy $\pi(a|x)$, the value function is defined as the expected return conditioned on a starting state $x$: $V^{\pi}(x)=\mathbb{E}_{\pi}\left[\sum_{t\geq 0}\gamma^{t}r_{t}|x_{0}=x\right]$, where $r_{t}=r(x_{t},a_{t})$ is the reward at time $t$. The value function is the unique stationary point of the Bellman operator $\mathcal{T}_{p}(V)(x)=\mathbb{E}_{\pi(a|s)}\left[r(x,a)+\gamma\mathbb{E}_{p(x^{\prime}|x,a)}\left[V(x^{\prime})\right]\right]$. The action-value function is defined similarly as $Q^{\pi}(x,a)=\mathbb{E}_{\pi}\left[\sum_{t\geq 0}\gamma^{t}r_{t}|x_{0}=x,a_{0}=a\right]$. In this work, we do not assume access to the reward functions and denote learned approximation as $\hat{r}(x,a)$.
Model-based RL:
An environment model is a function that approximates the behavior of the transition kernel of the ground-truth MDP. [^1] In this paper, we will generally use the term model to refer to an environment model, not to a neural network, to keep consistent with the reinforcement learning nomenclature. Learned models are used to augment the reinforcement learning process in several ways (Sutton, 1990; Janner et al., 2019; Hafner et al., 2020), we focus on the MVE (Buckman et al., 2018) and SVG (Amos et al., 2021) algorithms presented in details in Subsection 2.2. These were chosen as they fit the decision-aware model learning framework and have been used successfully in strong algorithms such as Dreamer (Hafner et al., 2021).
We will use $\hat{p}$ to refer to a learned probabilistic models, and use $\hat{f}$ for deterministic models. When a model is used to predict the next ground-truth observation $x^{\prime}$ from $x,a$ (such as the model used in MBPO (Janner et al., 2019)) we call it an observation-space models. An alternative are latent models of the form $\hat{p}(z^{\prime}|z,a)$, where $z\in\mathcal{Z}$ is a representation of a state $x\in\mathcal{X}$ given by $\phi:\mathcal{X}\rightarrow\mathcal{Z}$. In summary, observation-space models seek to directly predict next states in the representation given by the environment, while latent space models can be used to reconstruct learned features of the next state. We present further details on latent model learning in Subsection 3.1.
The notation $\hat{x}^{(n)}_{i}$ refers to an $n$-step rollout obtained from a model by starting in state $x_{i}$ and sampling the model for $n$-steps. We will use $\hat{p}^{n}\left(x^{(n)}\middle|x\right)$ and $\hat{f}^{n}(x)$ to refer to the $n$-th step model prediction.
@@ 2.1 Decision-Aware Model Losses
The losses of the decision-aware learning framework share the goal of finding models that provide good value function estimates. Instead of simply learning a model using maximum likelihood estimation, the losses are based on differences in value prediction. We present the loss functions below to highlight the close relationship as well as crucial differences between the two. We assume a standard dataset of state, action, and reward sequences $\mathcal{D}=\{x_{i_{1}},a_{i_{1}},r_{i_{1}},\dots,x_{i_{n}},a_{i_{n}},r_{i_{n}}\}_{i=1}^{N}$ with $x_{i_{1}}$ sampled from $\mu$ and subsequent states sampled according to the transition function. $N$ denotes the number of samples, and $n$ the sequence length. As all considered algorithms are off-policy, the actions can come from different past policies. Note that the index $1$ here denotes the start of a (sub-)sequence, not the start of an episode.
IterVAML (Farahmand, 2018):
To compute the IterVAML loss, value function estimates are computed over the model prediction and a real sample drawn from the training distribution. The loss minimizes the squared difference between them. While the original paper only considers a single-step variant, we present a multi-step version of the loss (for a discussion on the differences, see Subsection A.5). The resulting IterVAML loss is
^$\hat{\mathcal{L}}^{n}_{\text{IterVAML}}(\hat{p};\hat{V},\mathcal{D})$ $=\frac{1}{N\cdot n}\sum_{i=1}^{N}\sum_{j=1}^{n}\biggl{|}\underbrace{\mathbb{E}_{\hat{x}^{(j)}\sim\hat{p}^{j}(\cdot|x_{i_{1}},a_{i_{1}})}\left[\hat{V}\left(\hat{x}^{(j)}\right)\right]}_{\text{Model value estimate}}-\underbrace{\hat{V}\left(x_{i_{j}}\right)}_{\text{Ground truth value estimate}}\biggr{|}^{2}.$ (1)
MuZero (Schrittwieser et al., 2020):
MuZero proposes a similar loss. However, the next state value function over the real sample is estimated via an approximate application of the Bellman operator. This core difference between both losses is illustrated in Figure 1. Similar to Mnih et al. (2013), MuZero uses a target network for the bootstrap estimate, which we denote as $V_{\text{target}}$. The MuZero loss is introduced with using deterministic models, which is why we present it here as such. The extension to stochastic models is non-trivial, as we show in Subsection 4.2 and discuss further in Subsection A.2. The loss is written as
^$\hat{\mathcal{L}}^{n}_{\text{MuZero}}\left(\hat{f},\hat{V};\mathcal{D},V_{\text{target}}\right)$ $=\frac{1}{N\cdot n}\sum_{i=1}^{N}\sum_{j=i}^{n}\biggl{|}\underbrace{\hat{V}\left(\hat{f}^{j}(x_{i_{1}},a_{i_{1}})\right)}_{\text{Modelvalue estimate}}-\underbrace{\left[r_{i_{j}}+\gamma V_{\text{target}}\left(x_{i_{j+1}}\right)\right]}_{\text{Ground truth bootstrap value estimate}}\biggr{|}^{2}.$ (2)
Note that the real environment reward and next states are used in the $j$-th step. The MuZero loss is also used for updating both the value function and the model, while IterVAML is only designed to update the model and uses a model-based bootstrap target to update the value function.
Stabilizing loss:
@@
Decision-aware losses can be insufficient for stable learning  (Ye et al., 2021), especially in continuous state-action spaces (Hansen et al., 2022). Their performance is improved greatly by using a stabilizing or auxiliary loss. In practice, latent self-prediction, based on the Bootstrap your own latent (BYOL) loss (Grill et al., 2020), has shown to be a strong and simple option for this task. Several different versions of this loss have been proposed for RL (Gelada et al., 2019; Schwarzer et al., 2021; Tang et al., 2022), we consider a simple $n$-step variant:
@@
^$\hat{\mathcal{L}}^{n}_{\text{latent}}\left(\hat{f},\phi;\mathcal{D}\right)=\frac{1}{N\cdot n}\sum_{i=1}^{N}\sum_{j=1}^{n}\left[\hat{f}^{(j)}(\phi(x_{i_{1}}),a_{i_{1}})-\text{stop-grad}\left[\phi(x_{i_{j}})\right]\right]^{2}.$ (3)
Theoretical analysis of this loss by Gelada et al. (2019) and Tang et al. (2022) shows that the learned representations provide a good basis for value function learning.
2.2 Actor-critic learning
Value leaning:
Both IterVAML and MuZero can use the model to approximate the value function target. In its original formulation, MuZero is used with a Monte Carlo Tree Search procedure to expand the model which is not directly applicable in continuous state spaces. A simple alternative for obtaining a value function estimate from the model in continuous state spaces is known as model value expansion (MVE) (Feinberg et al., 2018). In MVE, short horizon roll-outs of model predictions $[\hat{x}^{(0)},\dots,\hat{x}^{(n)}]$ are computed using the current policy estimate, with a real sample from the environment as a starting point $x^{(0)}$. Using these, n-step value function targets are computed $Q_{\text{target}}^{j}\left(\hat{x}^{(j)},\pi\left(\hat{x}^{(j)}\right)\right)=\sum_{i=j}^{n-1}\gamma^{i-j}\hat{r}\left(\hat{x}^{(i)},\pi\left(\hat{x}^{(i)}\right)\right)+\gamma^{n-j}\bar{Q}\left(\hat{x}^{(n)},\pi\left(\hat{x}^{(n)}\right)\right)$. These targets represent a bootstrap estimate of the value function starting from the j-th step of the model rollout. For our IterVAML experiments, we used these targets together with TD3 (Fujimoto et al., 2018) to learn the value function, for MuZero, we used these targets for $V_{\text{target}}$ in Equation 2.
Figure 2: Comparison on the use of an explicit latent space with different loss functions. All losses improve in performance from the addition of an explicit latent space transformation. Dashed lines represent the mean of the latent space result.
Policy learning:
In continuous control tasks, estimating a policy is a crucial step for effective learning algorithms. To improve policy gradient estimation with a model, stochastic value gradients (SVG) (Heess et al., 2015; Amos et al., 2021) can be used. These are obtained by differentiating the full n-step MVE target with regard to the policy. As most common architectures are fully differentiable, this gradient can be computed using automatic differentiation, which greatly reduces the variance of the policy update compared to traditional policy gradients. We evaluate the usefulness of model-based policy gradients in Figure 6, with deterministic policy gradients (Silver et al., 2014; Lillicrap et al., 2016) as a model-free baseline.
3 Latent Decision Aware Models
Previous work in decision-aware MBRL (Farahmand et al., 2017; Farahmand, 2018; Schrittwieser et al., 2020; Grimm et al., 2021) uses differing model architectures, such as an ensemble of stochastic next state prediction networks in the case of Voelcker et al. (2022) or a deterministic latent model such as in Schrittwieser et al. (2020); Hansen et al. (2022). In this section, we highlight the importance of using latent networks for decision-aware losses. We call these variants of the investigate algorithms LAtent Model-Based Decision-Aware (LAMBDA) algorithms, or $\lambda$-IterVAML and $\lambda$-MuZero for short, to stress the importance of combining latent models with decision-aware algorithms. We focus on IterVAML and MuZero as the most prominent decision-aware algorithms.
3.1 The importance of latent spaces for applied decision-aware model learning
One of the major differences between MuZero and IterVAML is that the former uses a latent model design, while the IterVAML algorithm is presented without a specific architecture. When used with a state-space prediction model, IterVAML has been shown to diverge, leading to several works claiming that it is an impractical algorithm (Lovatto et al., 2020; Voelcker et al., 2022). However, there is no a priori reason why a latent model should not be used with IterVAML, or why MuZero has to be used with one. In prior work, sharp value function gradients have been hypothesized to be a major concern for IterVAML (Voelcker et al., 2022). The introduction of a learned representation space in the model through a latent embedding can help to alleviate this issue.
The full loss for each experiment is $\mathcal{L}^{n}_{\text{MuZero/IterVAML}}+\mathcal{L}^{n}_{\text{latent}}+\mathcal{L}^{n}_{\text{reward}}$, where $\mathcal{L}_{\text{reward}}$ is an MSE term between the predicted and ground truth rewards. As a baseline, we drop the decision-aware loss and simply consider BYOL alone. For all of our experiments we use eight random seeds and the shaded areas mark three $\sigma$ of the standard error of the mean over random seeds, which represents a $99.7\%$ certainty interval under a Gaussian assumption. Full pseudocode is found in Subsection C.2.
As is evident from Figure 2 the latent space is highly beneficial to achieve good performance. In higher dimensional experiments and without stabilizing losses, we find that IterVAML diverges completely (see Appendix D) This highlights that negative results on the efficacy of IterVAML are likely due to suboptimal choices in the model implementation, not due to limitations of the algorithm.